You need to dissolve CaCl2 in water to make a mixture that is 39.5% calcium chloride by mass. If the total mass of the mixture is 113.0 g, what masses of CaCl2 and water should be used? Mass of CaCl2 Mass of water
Answer #1

CaCl2 --> 39.5% by mass

total mass = 113 g

so..

39.5/100 * 113 = 44.635 g of CaCl2

so..

mass of water = 113-44.635 = 68.365 g of water